OpenStudy (itrymath):

A textbook company keeps track of its daily cost for printing whole textbooks to different schools but can only print up to 500 books a day. The daily cost (C) to print x number of textbooks is given by C(x) = 3.50x + 1200. Using the function, determine one possible daily cost by evaluating the function for any input of your choice. For what value of x does C(x) = 1900? Interpret the statement C(20) = 1270. Interpret the statement C(w) = p. Explain whether C(40.5) = 1341.75 is correct and possible given the context of this problem.

OpenStudy (itrymath):

@3mar

OpenStudy (itrymath):

@CamPayne

OpenStudy (campayne):

Well, The daily cost (C) to print x number of textbooks is given by : C(x) = 3.50x + 1200 1. C(x) = 1900 gives : C(x) = (7/2)x + 1200 = 1900 therefore : (7/2)x = = 1900 - 1200 = 700 finally : x = 2 . 700/7 = 200 books 2. C(20) = 1270 well, with the given function, first let's procced to the verification of the statement ... : c(20) = (7/2) . 20 + 1200 = 70 + 1200 = 1270 ---> OK means : the cost for printing 20 books is : $1270 3. Explain whether C(40.5) = 1341.75 is correct and possible given the context of this problem. same procedure : verify the statement ... 40.5 = 91/2 and C(40.5) = (7/2) . 91/2) + 1200 = 159.25 + 1200 = 1359.25 so C(40.5) = 1359.25 -----> the given cost $1341.75 is not correct

OpenStudy (3mar):

Sorry I was on other question!

OpenStudy (itrymath):

im confused

OpenStudy (itrymath):

@3mar

OpenStudy (3mar):

Well, I am here.

OpenStudy (itrymath):

how does he get 7/2

OpenStudy (3mar):

where?

OpenStudy (3mar):

3.5=7/2 Done!

OpenStudy (itrymath):

oooo

OpenStudy (3mar):

It happens to all of us!

OpenStudy (itrymath):

how does x = 2?

OpenStudy (itrymath):

C(x) = 3.50x + 1200 1. C(x) = 1900 gives : C(x) = (7/2)x + 1200 = 1900 therefore : (7/2)x = = 1900 - 1200 = 700 finally : x = 2 . 700/7 = 200 books

OpenStudy (3mar):

x doe not equal to 2 x is equal to 2×700÷7 Done!

OpenStudy (3mar):

got it? \[x=\frac{ 2*700 }{ 7 }=200\]

OpenStudy (itrymath):

oh because the way he wrote it didn't make sense thanks

OpenStudy (3mar):

Yes.

OpenStudy (3mar):

* . × all with the same meaning!

OpenStudy (itrymath):

i know

OpenStudy (3mar):

Good!

OpenStudy (3mar):

Any more questions?

OpenStudy (itrymath):

you guys can go help other people ill tag you if i need your help

OpenStudy (3mar):

Thank you for the medal!

OpenStudy (itrymath):

c(20) = (7/2) . 20 + 1200 = 70 + 1200 = 1270 ---> OK means : the cost for printing 20 books is : $1270